- 1 15. Let u= and v= 0 1. What is the dot product of u and v? -1] d) e) 2 L 1] b)-2 a) -1 c)0
Answer #1

ū= (2-1,0,1), V = 11,4,2;-) ū. = 9.6-1) + (-1).(4) + (0.).&#1.1-1) = + y + 0 = -2